OpenStudy (anonymous):

what is the equation of the parabola that opens to the left, has its vertex at the origin and passes through (-2,3)???

OpenStudy (anonymous):

not even close...

OpenStudy (anonymous):

If a parabola opens sideways, which variable is squared, x or y?

OpenStudy (anonymous):

y is squared

OpenStudy (anonymous):

How do you know?

OpenStudy (anonymous):

The squaring of the variables in the equation of the parabola determines where it opens: When y is squared and x is not, the axis of symmetry is horizontal and the parabola opens left or right.

OpenStudy (anonymous):

Right. Now, the vertex is at the origin, so we don't even have to subtract anything from either variable to translate it. So, our equation will look like\[y^{2} = cx\]How might we find c?

OpenStudy (anonymous):

this is where im lost

OpenStudy (anonymous):

What was the other piece of information they gave us in the problem?

OpenStudy (anonymous):

vertex passes through the origin

OpenStudy (anonymous):

I used that when I said we didn't have to subtract from x or y to translate the parabola. What have we not yet used?

OpenStudy (anonymous):

the points (-2,3)

OpenStudy (anonymous):

Right. How can we use that fact and y^2 = cx to find c?

OpenStudy (anonymous):

y^2=c(-2)

OpenStudy (anonymous):

Close. We know the point (-2, 3) will lie on the graph of our equation. That means (-2, 3) has to be a solution to our equation, y^2 = cx.

OpenStudy (anonymous):

so y^2 =3^2=9 2/9y?

OpenStudy (anonymous):

no, -(9/2)y

OpenStudy (anonymous):

x...

OpenStudy (anonymous):

Remember, the point (-2, 3) means x=-2 and y=3. We can plug both of those numbers in to our equation. y^2 = cx 3^2 = c(-2) c = -(9/2) So putting it all together, what will the final equation be?

OpenStudy (anonymous):

y^2=-(9/2)x

OpenStudy (anonymous):

Excellent!

OpenStudy (anonymous):

thanks for the help!